OpenStudy (anonymous):

whats z^2z^4/z^3Z^-1

OpenStudy (anonymous):

\[z ^{2} z ^{4}/z ^{3}z ^{-1}\]

OpenStudy (dinamix):

\[z^2z^4/z^3z ^{-1}= z^1/z^3 = 1/z^2\]

OpenStudy (dinamix):

u study number complexe

OpenStudy (anonymous):

how did you get z^1 on the top and z^3 on the bottom?

OpenStudy (anonymous):

@dinamix

OpenStudy (anonymous):

isn't it 2+4 which is 6 and the 3+ (-1)=2

OpenStudy (dinamix):

z^4/z^3 = Z^(4-3) = Z^1

OpenStudy (anonymous):

how're you getting those numbers though

OpenStudy (dinamix):

and z^2/z^-1 = z^(2-(-1)) =z^3

OpenStudy (anonymous):

ohhh okay i see what you did there! this problem was so easy but it wars confusing for second thanks for your help

OpenStudy (dinamix):

its when u see Basis of divided like this u have make difference

OpenStudy (dinamix):

\[z^x * z^y = z ^{x+y}\]

OpenStudy (dinamix):

\[z^x/z^y = z^{x-y}\]

OpenStudy (dinamix):

@Anzhbar